# Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3

### v1

@@ Abstract
As the parameters of large language models (LLMs) expand, the computational cost of fine-tuning the entire model becomes prohibitive. To address this challenge, we introduce a parameter-efficient fine-tuning (PEFT) method, Principal Singular values and Singular vectors Adaptation (PiSSA), which optimizes a significantly reduced parameter space while achieving or surpassing the performance of full-parameter fine-tuning. PiSSA is inspired by Intrinsic SAID, which suggests that pre-trained, over-parametrized models inhabit a space of low intrinsic dimension. Consequently, PiSSA represents a matrix $W\in\mathbb{R}^{m\times n}$ within the model by the product of two trainable matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, plus a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ for error correction. Singular value decomposition (SVD) is employed to factorize $W$, and the principal singular values and vectors of $W$ are utilized to initialize $A$ and $B$. The residual singular values and vectors initialize the residual matrix $W^{res}$, which keeps frozen during fine-tuning. Notably, PiSSA shares the same architecture with Low-Rank Adaptation (LoRA), which hypothesizes that changes in model parameters $\Delta W$ form a low-rank matrix. However, LoRA approximates $\Delta W$ through the product of two matrices, $A$, initialized with Gaussian noise, and $B$, initialized with zeros, while PiSSA initializes $A$ and $B$ with principal singular values and singular vectors of the original matrix $W$. Given that the principal singular values and vectors capture the essence of a low-rank matrix, PiSSA can better approximate the outcomes of full-parameter fine-tuning at the beginning by changing the essential parts while freezing the “noisy” parts. In comparison, LoRA freezes the original matrix and updates the “noise”. This distinction enables PiSSA to convergence much faster than LoRA and also achieve better performance in the end. On five common benchmarks, PiSSA outperforms LoRA on all of them using exactly the same setups except for a different initialization. On GSM8K, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, outperforming LoRA’s 67.7% by 5.16%. Due to the same architecture, PiSSA inherits many of LoRA’s advantages, such as parameter efficiency and compatibility with quantization. Leveraging a fast SVD method, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
@@ 1 Introduction
A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix, instead optimizing the injected, significantly smaller low-rank matrices. Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times. Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks. By integrating the quantization of pre-trained models, LoRA also enables reducing the average memory requirements by 16 times [10]. Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full precision base models.
@@
Different from LoRA, a significantly less important part $W^{res}$ is kept frozen, while the essential part $W^{pri}$ is fully tunable from the beginning, which enables PiSSA to fit the training data faster and better.
@@ 2 Related Works
Differing from LoRA’s approach of learning low-rank approximations of weight updates, our PiSSA method innovates by utilizing the low intrinsic dimension of pre-trained weight directly. It initializes updatable adapter layers with principal singular values and vectors from the model’s weights, preserving the essential capabilities while enabling effective fine-tuning. This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like Low-rank Adjustment for enhanced fine-tuning efficiency.
@@ 3 Method
Given $W=W^{res}+AB$, the introduction of PiSSA in the beginning of training does not compromise the model’s capabilities. Considering $S_{[r:]}\approx 0$, thus rendering $W^{res}\approx 0$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities. Then, the fine-tuning of $AB$ mirrors the process of fine-tuning the entire model, despite using much less parameters. In contrast, LoRA initializes the adapter with Gaussian noise and zeros, thus the principal part of $W$ is not updatable in the early stage of the fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include but not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, the quantization of the residual model to decrease memory consumption during forward propagation in training, and easy deployment. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\{A,B\}$ instead. When we want to utilize a fine-tuned PiSSA, similar to initializing one, we first perform singular value decomposition on the base model (which is very fast as will be shown in our experiments) and then assign $\{A,B\}$ to the adapter. Finally, one pre-trained model can accommodate multiple $\{A,B\}$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.
@@ 4 Experiments
We conduct comparative experiments between PiSSA, LoRA, and full parameter fine-tuning on LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50]. These comparing cover a wide range of tasks, including math [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6]. The findings reveal that fine-tuning adapters initialized with principal singular values and vectors (PiSSA) outperform those initialized with intermediate or less significant ones. This evidences that direct fine-tuning of the model’s principal components leads to superior outcomes (Section 4.1). PiSSA demonstrates superior performance on downstream tasks (Section 4.3). Subsequently, we conduct experiments across various ranks, observing that PiSSA converges more swiftly and aligns more closely with the training data compared to LoRA, exhibiting robust superiority under the same trainable parameter configurations (Section 4.2). Finally, we employ subsequent experiments to demonstrate that the Fast SVD [56] technique helps our PiSSA achieve a balance between initialization speed and performance (Section 4.4).
@@ 4.1 Verifying the importance of principal singular values and vectors
4.2 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
(b) Final training loss across different ranks.

### v2

@@ Abstract
As the parameters of large language models (LLMs) expand, the computational cost of fine-tuning the entire model becomes prohibitive. To address this challenge, we introduce a parameter-efficient fine-tuning (PEFT) method, Principal Singular values and Singular vectors Adaptation (PiSSA), which optimizes a significantly reduced parameter space while achieving or surpassing the performance of full-parameter fine-tuning. PiSSA is inspired by Intrinsic SAID, which suggests that pre-trained, over-parametrized models inhabit a space of low intrinsic dimension. Consequently, PiSSA represents a matrix $W\in\mathbb{R}^{m\times n}$ within the model by the product of two trainable matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, plus a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ for error correction. Singular value decomposition (SVD) is employed to factorize $W$, and the principal singular values and vectors of $W$ are utilized to initialize $A$ and $B$. The residual singular values and vectors initialize the residual matrix $W^{res}$, which keeps frozen during fine-tuning. Notably, PiSSA shares the same architecture with Low-Rank Adaptation (LoRA), which hypothesizes that changes in model parameters $\Delta W$ form a low-rank matrix. However, LoRA approximates $\Delta W$ through the product of two matrices, $A$, initialized with Gaussian noise, and $B$, initialized with zeros, while PiSSA initializes $A$ and $B$ with principal singular values and singular vectors of the original matrix $W$. Given that the principal singular values and vectors capture the essence of a low-rank matrix, PiSSA can better approximate the outcomes of full-parameter fine-tuning at the beginning by changing the essential parts while freezing the “noisy” parts. In comparison, LoRA freezes the original matrix and updates the “noise”. This distinction enables PiSSA to convergence much faster than LoRA and also achieve better performance in the end. On five common benchmarks, PiSSA outperforms LoRA on all of them using exactly the same setups except for a different initialization. On GSM8K, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, outperforming LoRA’s 67.7% by 5.16%. Due to the same architecture, PiSSA inherits many of LoRA’s advantages, such as parameter efficiency and compatibility with quantization. Furthermore, PiSSA reduces the 4-bit quantization error in LLaMA 2-7B by 18.97%, resulting in a substantial improvement in fine-tuning performance. On the GSM8K benchmark, PiSSA achieves an accuracy of 49.13%, surpassing the performances of QLoRA at 39.8% and LoftQ at 40.71%. Leveraging a fast SVD technique, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
@@ 1 Introduction
A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix, instead optimizing the injected, significantly smaller low-rank matrices. Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times. Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks. By integrating the quantization of pre-trained models, LoRA also enables reducing the average memory requirements by 16 times [10]. Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full-precision base models.
@@
Different from LoRA, a significantly less important part $W^{res}$ is kept frozen, while the essential part $W^{pri}$ is fully tunable from the beginning, which enables PiSSA to fit the training data faster and better. Since the high-frequency principal component $W^{pri}$ is preserved in the adapter, when fine-tuning PiSSA incorporated with quantization, PiSSA significantly reduced quantization errors in the initial stages compared to using QLoRA.
@@ 2 Related Works
Differing from LoRA’s approach of learning low-rank approximations of weight updates, our PiSSA method innovates by utilizing the low intrinsic dimension of pre-trained weight directly. It initializes updatable adapter layers with principal singular values and vectors from the model’s weights, preserving the essential capabilities while enabling effective fine-tuning. This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like quantization the residual model for reducing computation cost.
@@ 3 Method
Given $W=W^{res}+AB$, the introduction of PiSSA in the beginning of training does not compromise the model’s capabilities. Considering $S_{[:r]}\gg S_{[r:]}$, thus rendering $W^{pri}\gg W^{res}$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities. Then, the fine-tuning of $AB$ mirrors the process of fine-tuning the entire model, despite using much less parameters. In contrast, LoRA initializes the adapter with Gaussian noise and zeros, thus the principal part of $W$ is not updatable in the early stage of the fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include, but are not limited to, the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment, during which process PiSSA can significantly reduce the initial quantization error in the initial stages. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\Delta AB$ instead. As shown in Appendix A, leveraging solely the $\Delta AB$ facilitates their seamless integration with the original pre-trained models. Finally, one pre-trained model can accommodate multiple $\Delta AB$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.
@@ 4 Experiments
We conducted comparative experiments among PiSSA, LoRA, and full parameter fine-tuning on three models: LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50]. Our comparison spanned a diverse range of tasks, encompassing mathematics [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6]. PiSSA exhibited superior performance across these downstream tasks (Section 4.1). When combined with quantization techniques, PiSSA effectively reduces quantization errors. Compared to QLoRA [10] and LoftQ [47], PiSSA exhibits superior performance both during initialization and fine-tuning stages (Section 4.2). The ablation studies, which varied across different ranks, showed that PiSSA converged more swiftly and aligned more closely with the training data than LoRA, demonstrating robust superiority under equivalent trainable parameter configurations (Section 4.3). Further experiments demonstrated that employing the Fast SVD technique [56] allowed PiSSA to achieve a balance between initialization speed and performance (Section 4.4). Lastly, adapters initialized with principal singular values and vectors (PiSSA) were compared to those initialized with intermediate or less significant ones, revealing that direct fine-tuning of the model’s principal components yielded better outcomes (Section 4.5).
@@ 4.1 Comparative Analysis of PiSSA, LoRA, and Full Fine-Tuning Across Various Tasks
4.2 Comparative Analysis of QLoRA, LoftQ, and PiSSA in Quantization Fine-tuning
In this section, we compare the performance of QLoRA, LoftQ, and PiSSA in fine-tuning with quantization. All base models are quantized to ‘nf4’ precision and frozen, while the adapters are trained using ‘bfloat16’ precision. We begin by comparing the quantization errors of the three methods. For each method, the quantized baseline model is combined with its adapter, and the difference between the original weights and this combination serves as the error matrix. We measure the magnitude of the errors by the average nuclear norm of all error matrices. Using the quantization error of the original model without any adapter as a baseline, the errors for QLoRA, LoftQ, and PiSSA are divided by this baseline to express them as error rates. As QLoRA’s adapter employs Gaussian-zero initialization, the error rate does not decrease; LoftQ can eliminate quantization errors through multiple singular value decompositions, but for consistency with PiSSA, only one decomposition is performed. PiSSA requires a full-precision singular value decomposition, storing the principal singular values and vectors in the adapter before quantizing the residual model. We train LLaMA 2-7B on MetaMathQA and validate the effectiveness of the three methods on GSM8K and MATH datasets. The reductions in error and the training results are displayed in Figure 2.
Figure 2: Comparing the quantization error, the fine-tuning loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets.
The results, shown in Figure 2, reveal that PiSSA significantly outperforms LoftQ in reducing errors, achieving a 19% reduction at rank 128, as opposed to LoftQ’s 8%. LoftQ’s fine-tuning improvements over QLoRA are not significant, and in certain settings, it even underperforms compared to QLoRA. PiSSA effectively reduces quantization errors, and the training of principal singular values and vectors additionally provides benefits, thus substantially enhancing training outcomes in comparison to QLoRA and LoftQ.
4.3 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
(b) Final training loss across different ranks.
4.4 Fast Singular Value Decomposition
@@ 5 Conclusion
Appendix A Equivalently Convert PiSSA into LoRA
The advantage of PiSSA lies in its ability to significantly enhance training outcomes during the fine-tuning phase. After training, it allows for the direct sharing of the trained matrices $A$ and $B$. However, users need to perform singular value decomposition on the original model, which requires additional time. When employing fast singular value decomposition, there can be slight inaccuracies. More importantly, utilizing PiSSA for inference necessitates altering the parameters of the original model, which can be inconvenient when using multiple adapters, especially when some adapters might be disabled or activated. Therefore, we recommend converting the trained PiSSA module equivalently into a LoRA module, thereby eliminating the need to modify the original model’s parameters during sharing and usage. In the initialization phase, PiSSA decomposes the original matrix into principal components and a residual matrix: $W=W^{res}+A_{0}\times B_{0}$. Upon completion of training, the model adjusts the weights as follows: $W+\Delta W=W^{res}+A\times B$. Thus, the modification of the model weights by PiSSA is given by:
$\Delta W$ $=A\times B-A_{0}\times B_{0}$ (6)
$=[A\,|\,A_{0}]\times\begin{bmatrix}B\\ -B_{0}\end{bmatrix}$ (7)
$=A^{{}^{\prime}}B^{{}^{\prime}}.$ (8)
[…]
Appendix B Training Loss under Various Ranks
(a) Variation of loss during the fine-tuning of LLaMA 2-7B with PiSSA throughout the training phase.
(b) Variation of loss during the fine-tuning of Mistral-7B-v0.1 with PiSSA throughout the training phase.
(c) Variation of loss during the fine-tuning of Gemma-7B with PiSSA throughout the training phase.
Figure 5: The change in loss during the training process with respect to ranks [2, 4, 8, 16, 32, 64, 128].
Appendix C Quantization Error for More Layers
Figure 6: Comparison of quantization errors in QLoRA, LoftQ, and PiSSA across k_proj, v_proj, o_proj and gate_proj, up_proj, down_proj layers.